OpenStudy (cutiecomittee123):

Solve cos (x) (cos x +1) = 0 I dont even know where to start here honestly

OpenStudy (cutiecomittee123):

@campbell_st can you help me? please

OpenStudy (campbell_st):

ok... so this is a quadratic function if wrote it as x(x + 1) = 0 could you solve it..?

OpenStudy (cutiecomittee123):

Yes I can so basically just distribute and then retricelike terms and solve?

OpenStudy (campbell_st):

but the basics are that the equation is already factored... so if either factor is zero, the who equation is zero... so you need to solve cos(x) = 0 and cos(x) + 1 = 0 those angles will be the solutions...

OpenStudy (cutiecomittee123):

So cos x = -1 then I just find where this is true on the unit cirlce and use the cooresponding radians right?

OpenStudy (cutiecomittee123):

but then i only get pi as an answer

OpenStudy (cutiecomittee123):

oh but when it equals zero is important to so pi and 3pi/2 and pi/2

OpenStudy (campbell_st):

yes use the unit circle, but I find the curve y = cos(x) a little easier  I don't know if you need to answer in radians or degrees hope it helps